# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ 3.2.2 The modern view on learning
The logic of inductive bias is sound. On a deep level, there really is no free lunch in learning, even deep learning (Mitchell,, 1980; Adam et al.,, 2019). However, developments in deep learning have forced massive revisions to conventional ways of thinking about inductive bias.
@@
The conventional wisdom was that a learning model must be restricted lest it overfit (Bishop,, 2006, §1.1). That is, an unrestricted model may memorize the training data, or find some other pathological solution which lurks in the depths of its vast hypothesis space, as in the red hypotheses in Figure 3.2.1. By restricting the hypothesis space, one provides strong inductive bias, so that pathological solutions can be ruled out and good generalization can be guaranteed. The logic is the same as Universal Grammar: explaining generalization requires that we restrict the set of hypotheses (grammars) that a model (learner) may entertain.
@@
However, the empirical finding from deep learning is that overparameterized networks, which are far more powerful than would be needed to memorize their training data, do generalize, often better than restricted ones (Belkin et al.,, 2019; Zhang et al.,, 2021). In machine learning, overfitting has typically been measured in terms of performance on a held-out dataset, where one typically sees a U-shape curve as in Figure 2A: as models become more powerful in terms of the hypotheses they can express, at first they generalize to the held-out data poorly (underfitting), then they generalize well (the sweet spot), then after some point the more powerful models learn poorer hypotheses (overfitting). And yet in the late 2010s it was discovered that if you keep adding model power, the U-shape curve starts to descend again, in a phenomenon called double descent, shown in Figure 2B, indicating that learners start to find good generalizations after being able to memorize their training data. [^8] In fact, double descent had already been discovered in the statistical physics literature (Vallet et al.,, 1989, p. 319), but its significance for machine learning was not appreciated at the time. It seems that what appeared to be an insurmountable barrier turned out to actually be a kind of phase transition in how learning works: below a certain level of model capacity, one sees the usual picture of over- and under-fitting, but above that level, one sees a new kind of generalization emerging, where more powerful learners have superior performance in learning the right generalizations.
under-fittingover-fittingsweet spotTest riskTraining riskModel capacityRiskAunder-parameterizedover-parameterizedinterpolation thresholdTest riskTraining riskModel capacityRiskB
@@
Figure 2: Conceptual risk (probability of error) on training set and test set as a function of model capacity (the variety of hypotheses that the model can express). A. The classical picture from statistical learning theory. In the under-fitting regime, increasing model capacity reduces training and test risk, up to a point where one enters the over-fitting regime, in which increasing model capacity causes a decrease in training risk but an increase in test risk. This overfitting phenomenon is intuitively due to the model gaining the capacity to memorize the training set without forming any generalizations that would be useful on the test set. The goal of model fitting in this view is to find the sweet spot that minimizes training risk, often accomplished through methods such as deliberately reducing model capacity or early stopping. B. The modern picture arising empirically from deep learning. Classical statistical learning theory takes place in the under-parameterized regime. As model capacity increases beyond the point where the model can memorize the training data (the interpolation threshold), we enter the over-parameterized regime, where test risk decreases. Figure based on Belkin et al., (2019).
@@
These findings came as a surprise to statistical learning theorists (Belkin et al.,, 2019; Zhang et al.,, 2021) and have triggered an ongoing field-wide effort to rethink learning theory (for example, Poggio et al., 2020a, ; Martin et al.,, 2021; Martin and Mahoney,, 2021; Kuzborskij et al.,, 2021; Henighan et al.,, 2023; Attias et al.,, 2024). We believe these results should also make us rethink language learning and the role of restrictive formalisms in linguistic explanation. It is simply not the case that proper generalization requires restricted learners, nor even that there is a correlation between restrictedness and proper generalization. All the successes of modern machine learning are in contradiction to the view that effective learners need to operate in a restricted hypothesis space.
@@
But how can this be, if the idea of inductive bias was true? How did the conventional logic go wrong in practice? The key mistake was the conflation of model power with inductive bias, that is, the idea that inductive bias has to come from restrictions on the learner’s hypothesis space (Hubinger,, 2019). In fact, the modern empirical finding is that less restricted learners have stronger biases toward ‘simple’ hypotheses (Huh et al.,, 2024). Furthermore, the source of this simplicity bias in neural networks and related systems is not necessarily related to the hard limits of their expressivity. This is an area under active investigation—neural systems seem to be biased towards functions that are nearly linear or simple in other ways (Valle-Perez et al.,, 2018; Hahn et al., 2021b, ), likely as a result of the dynamics of gradient descent on the loss landscape induced by the model (Poggio et al., 2020b, ; Pezeshki et al.,, 2020; Merrill et al., 2021b, ; Hahn and Rofin,, 2024).
@@
In machine learning, these discoveries have catalyzed a change in focus, away from models whose architecture and representations are tailor-fit to their domain, and toward models that learn quickly in relatively unrestricted hypothesis spaces (Sutton,, 2019). The zeitgeist is now strongly against ‘building in’ domain knowledge or targeted biases, and rather focusing on finding powerful general learners. New model architectures are now routinely evaluated based on the slope of their error rate as a function of amount of training, called scaling laws (Kaplan et al.,, 2020), rather than based on the expressivity of their architecture.
@@ 3.2.3 The upshot for linguistic theory
This revolution in learning theory means that the research program of explaining language universals through restricted learners is deprived of its apparent logical inevitability. Previously, the logic was: learners must be formally restricted, and we see that human languages do have universal patterns, so maybe these necessary restrictions create the universal patterns. Nowadays, if this approach to linguistic explanation is to be maintained, the logic must be: learners are restricted even though this is not necessary and may be harmful for learning, and these restrictions create language universals. It is still a viable hypothesis, but it no longer seems so elegant. There is no logical necessity for a restrictive formalism to explain language learning, so the approach of “explanation by constrained description” no longer seems so explanatory. Furthermore, even if evolutionarily-contingent innate inductive biases in humans really are the key to the structure of human language, these inductive biases might not be expressible in terms of a discrete formalism or a limited hypothesis space for learners.
It is still the case that, whatever language models learn about language, they do it on the basis of orders of magnitude more linguistic data than human children are exposed to (Yedetore et al.,, 2023; Warstadt et al.,, 2023). While it has been suggested that formal linguistic structure, perhaps built-in to the architecture of the model, would provide the key inductive bias to make learning from small amounts of data possible, there is no evidence that this is the only missing ingredient, or even an important one, to close the gap between models and human children. [^9] Indeed, the analysis by Mollica and Piantadosi, (2019) suggests that hierarchical syntactic structure makes up only a very small proportion of the information necessary to learn a language. It is entirely possible that the key to faster and more human-like learning is not particular built-in formal constraints, but rather more expressive architectures (such as Kolmogorov–Arnold Networks: Liu et al.,, 2024), or biases towards domain-general compositional symbolic reasoning (McCoy and Griffiths,, 2023; Yang and Piantadosi,, 2022), or simply a different training regime (Murty et al.,, 2023), or the incorporation of multimodal data which provides rich side information about the structure of the environment that is being described in language (Wang et al.,, 2023). Attempts to inject linguistic structure into neural models in various ways do succeed in promoting the kinds of structural generalizations that humans find intuitive (Dyer et al.,, 2016; Futrell et al., 2019b, ; Wilcox et al., 2019b, ; Kim et al.,, 2019), and they do succeed in making language learning from data somewhat more sample efficient (Papadimitriou and Jurafsky,, 2020, 2023), but gains in terms of absolute overall performance are small, and so such models have not caught on. Nor is it the case that formal linguists have presented an alternative model with anything like the demonstrated practical language-learning capabilities of neural models.
@@
This is not to say that inductive biases are no longer important, nor that investigating inductive biases of humans and neural networks is not important: we will have much more to say on this in Section 4.4. Far from demoting inductive bias as a concern, language models open up the kinds of (possibly innate) inductive biases that we might look for in humans. The main upshot for linguistic theory is not to demote the importance of the learning problem, but to reshape how we think about it: away from reliance only on restrictive formalisms, and toward a broader landscape of hypotheses.
@@ 3.3 Language models and linguistic traditions
While language models may seem to involve concepts that are foreign to linguistics, they have in fact emerged in part from certain intellectual traditions of the study of language—although these traditions are not necessarily the ones that have been dominant in linguistics departments. Below we discuss the relationship between schools of linguistics and the development of language models and related language technologies.
In doing so, we focus primarily on linguistic structure, particularly related to syntax. This mirrors the historical structure of the field, and it also reflects the fact that language models make profound contributions to inquiry in these domains. But we note that linguistics also encompasses a rich tradition of the study of sociocultural aspects of language (Bucholtz and Hall,, 2005; Meyerhoff,, 2006), and that this work is increasingly important in AI. Insofar as LMs are also now (sometimes unwelcome) interlocutors in our linguistic communities due to the proliferation of widely used chatbots and the influx of automatically generated text into daily life, linguists are well-positioned to study the social and cultural role of LMs in linguistic communities.
@@ 3.3.1 The generative tradition of linguistics
A historical parallel is informative. In the early 19th century, while Newtonian mechanics provided a strong basis for certain parts of mechanical engineering, it did not on its own seem to provide a theory that answered pressing questions about the increasingly complex machines emerging from the Industrial Revolution, in particular (1) the principles underlying steam engines, and (2) the fact that perpetual motion machines appeared to be impossible. Carnot, (1803) and then Carnot, (1824) developed a theory of these phenomena in terms of a new ad-hoc concept of ‘moment of activity’ which was dissipated inexorably by irreversible shocks to the system—irreversible at the coarse-grained level of realistic observations and average quantities, but not at the level of fine-grained mechanics which had been the focus of theoretical physics. The ‘moment of activity’ eventually developed into the idea of entropy (Clausius,, 1865) and the whole field of thermodynamics followed.
@@
Entropy had to be discovered through engineering because the focus of purely theoretical physicists was on understanding fundamental dynamics and forces, and on developing mathematical methods of increasing elegance and sophistication. Explaining the behavior of complex machines—which behavior certainly must reduce to more basic mechanics anyway—was not on the agenda. And yet from the apparently ad-hoc project of understanding steam engines emerged a family of concepts that some theoretical physicists now view as more fundamental than even spacetime itself (Wheeler,, 1989; Cao et al.,, 2017).
@@
Turning back to linguistics: the reason that generative linguistic theory did not turn out to be helpful along the path toward language models was due to narrowness in focus. Just as the focus of theoretical physics in the early 19th century did not lead it to consider effective theories of steam engines, the focus of generative linguistics in the 20th and early 21st centuries did not lead it to consider complex systems like language models.
@@
Language models are, at bottom, models of the stream of language that is produced and comprehended by humans. The relevant theory is the theory of language use, production, comprehension, learning, and of human cognition more generally. Generative linguistics, on the other hand, has not focused on how to build a machine that produces and comprehends language, but rather how to build a language, conceived of as an abstract mental mapping between meaning (that is, a logical form or conceptual–intensional representation) and form (that is, a phonological form or sensorimotor representation) (Chomsky,, 2005). [^10] Early generative linguistics went further in its goals, claiming that the structures identified through generative analysis would not only provide a characterization of language, but also be a necessary and explicit part of any theory of how language could be learned and processed (Chomsky,, 1965, 1971; Bresnan,, 1982). See Stabler, (1983) for discussion of what he sees as confusion within the field as to whether or not generative linguistic theories are intended to be theories of the representations used by the brain during processing, as opposed to theories that constrain possible language. He concludes that Chomsky and others often conflated theories of grammars and theories of mental representation and processing, but that this position was not well grounded and that “it seems unlikely that linguists and psychologists really want to claim any such thing”. Furthermore, it was claimed that this kind of analysis must take center stage in the science of language, preceding any analysis of more complex systems for language processing, use, or learning (Chomsky,, 1965, Ch. 1).
@@
Given this focus, generative linguistics has had a huge influence on the engineering field of programming language design. Indeed, the basics of all modern programming languages are based on formal linguistic theory, incorporating explicit (often context-free) grammars and parsers, as a way of linking a stream of text to a representation of a computation. Programming languages are, perhaps, exactly the ‘esoteric dialects’ spoken by ‘elaborate machines’ in the Nabokov quote that starts this section.
@@
While this narrowness of focus was useful on its own terms, we believe it has limits. There are potentially fundamental insights to be gained now from the analysis of messy, complex, practical systems, just as happened in physics, and as is now happening in linguistics.

### v2

@@ 3.2.2 The modern view on learning
The logic of inductive bias is sound. On a deep level, there really is no free lunch in learning, even deep learning (Mitchell,, 1980; Wolpert et al.,, 1995; Baxter,, 2000; Adam et al.,, 2019). Asking how much one can learn ‘from the data alone’ without inductive bias is like asking how close a runner can get to the finish line without saying where they started. It’s just not a sensible question. However, developments in deep learning have forced revisions to conventional ways of thinking about how inductive bias arises.
@@
The conventional wisdom was that a learning model must be restricted in terms of the hypotheses it can consider, so that it does not overfit (Bishop,, 2006, §1.1). A relatively unrestricted model may memorize the training data, or find some other pathological solution which lurks in the depths of its vast hypothesis space, as in the red hypotheses in Figure 3.2.1. By restricting the hypothesis space, one provides strong inductive bias, so that pathological solutions can be ruled out and good generalization can be guaranteed. The logic is the same as Universal Grammar: explaining generalization requires that we restrict the set of hypotheses (grammars) that a model (learner) may entertain.
@@
However, the empirical finding from deep learning is that overparameterized networks, which are more than flexible enough to memorize their training data, do generalize, often better than relatively restricted ones (Belkin et al.,, 2019; Zhang et al.,, 2021). In machine learning, overfitting has typically been measured in terms of performance on a held-out dataset, where one typically sees a U-shape curve as in the left part of Figure 2: as models become more flexible in terms of the hypotheses they can express, at first they generalize to the held-out data poorly (underfitting), then they generalize well (the sweet spot), then after some point the more powerful models learn poorer hypotheses (overfitting). And yet in the late 2010s it was discovered that if you keep adding model power, the U-shape curve starts to descend again, in a phenomenon called double descent, shown in Figure 2. This pattern suggests that learners continue to find good generalizations even after being flexible enough to memorize their training data. [^11] In certain settings and with the right regularization, the best performance comes from models whose capacity is poised right at the point where training data and model parameters are balanced—the right regularization in these settings avoids the sharp spike in loss characteristic of double descent (Maloney et al.,, 2022, §4.2).
It seems that what appeared to be an insurmountable barrier turned out to be a kind of phase transition in how learning works. What looks like overfitting is, counterintuitively, ameliorated by adding even more parameters. Thus, one sees a new kind of generalization emerging, where less restricted learners have superior performance in learning the right generalizations.
@@
Figure 2: Expected error of a learning model on training set and test set as a function of model expressivity (the variety of hypotheses that the model can express). To the left of the blue line, we have the classical picture from statistical learning theory. In the underfitting regime, increasing model capacity reduces training and test error, up to a point where one enters the overfitting regime, in which increasing model capacity causes a decrease in error on the training data but an increase in error on held-out data. This overfitting phenomenon is intuitively due to the model gaining the capacity to memorize the training set without forming generalizations that would be useful on the test set. The goal of model fitting in this view is to find the sweet spot that minimizes test error, often accomplished through methods such as deliberately reducing model capacity or early stopping. However, modern deep learning has revealed that as model capacity increases beyond the point where the model can memorize the training data (the interpolation threshold, in blue), we enter a new overparameterized regime, where test error decreases, due to soft simplicity biases in the learner. Figure based on Belkin et al., (2019).
@@
These findings came as a surprise to statistical learning theorists (Belkin et al.,, 2019; Zhang et al.,, 2021) and have triggered an ongoing field-wide effort to rethink learning theory or to show how these unexpected findings are compatible with existing theory (for example, Poggio et al., 2020a, ; Martin et al.,, 2021; Martin and Mahoney,, 2021; Kuzborskij et al.,, 2021; Henighan et al.,, 2023; Attias et al.,, 2024). These results should also make us rethink language learning and the role of restrictive formalisms in linguistic explanation. It is simply not the case that proper generalization can only come from learners who are sharply restricted to small hypothesis spaces, nor even that there is a correlation between restrictedness and proper generalization. All the successes of modern machine learning are in contradiction to the view that generalization must come from hard restrictions on the hypothesis space.
@@
But how can this be, if the idea of inductive bias is right? How did the conventional logic go wrong in practice? The key mistake was the conflation of model power with inductive bias (Hubinger,, 2019). The modern empirical finding is that more flexible learners have stronger biases toward “simple” hypotheses (Huh et al.,, 2024): that is, although they do not restrict the range of hypotheses that can be considered, they do impose a soft notion of simplicity on those hypotheses. The logic is illustrated in Figure 3, following Wilson, (2025). Furthermore, the source of this simplicity bias in neural networks and related systems is not necessarily related to the hard limits of their expressivity. The simplicity bias in neural networks seems to be toward functions that are nearly linear or simple in other ways (Valle-Perez et al.,, 2018; Hahn et al., 2021b, ), likely as a result of the dynamics of gradient descent on the loss landscape induced by the model (Poggio et al., 2020b, ; Pezeshki et al.,, 2020; Merrill et al.,, 2021; Hahn and Rofin,, 2024).
Figure 3: Illustration of the logic of a soft inductive bias within a flexible hypothesis space, adapted from Wilson, (2025). Figures show a hypothesis space for a learner, including regions of bad generalization (overfitting) and good generalization. The learner starts at the dot and, throughout learning, moves in the direction indicated by the arrow. A. Given a uniform bias over a highly flexible, relatively unrestricted hypothesis space, the learner is likely to overfit and form pathological generalizations. B. One strategy to impart useful inductive bias to a learner is to restrict the hypothesis space. However, this may miss the good generalizations. C. A better strategy, which is part of what underlies neural network success, is to keep a highly flexible, relatively unrestricted hypothesis space, but impart a soft simplicity bias which guides the learner towards the good generalizations.
@@
In machine learning, these discoveries have catalyzed a change in focus, away from models whose architecture and representations are tailor-fit to their domain, and toward models that learn quickly in relatively unrestricted hypothesis spaces (Sutton,, 2019). The zeitgeist is now strongly against using domain knowledge to restrict the behavior of learners, even when researchers have a strong sense of what the relevant domain knowledge is.
@@ 3.2.3 The upshot for linguistic theory
These surprising results from deep learning mean that the research program of achieving explanatory adequacy through restricted learners is deprived of its apparent logical inevitability. Previously, the logic was: learners must be formally restricted, and we see that human languages do have universal patterns, so maybe these necessary restrictions create the universal patterns. Nowadays, if this approach to linguistic explanation is to be maintained, the logic must be: learners have hard formal restrictions even though this is not necessary and may be harmful for learning, and these restrictions create language universals. It is still a viable hypothesis, but it loses its elegance.
Furthermore, even if language-specific innate inductive biases in humans really are the key to the structure of human language, these inductive biases might not be expressible in terms of a categorical symbolic formalism or a sharply limited hypothesis space for learners. Inductive biases in modern neural models are soft and seem to arise from a complex interplay of training dynamics, objective function, and model architecture, with the hard limits of model expressivity playing a relatively minor role.
Overall, “explanation by constrained description” no longer seems so explanatory, at least not without further stipulations. Language learning and language universals may well be better captured by a highly flexible, less constrained formalism for linguistic representation—one which on its own could capture non-linguistic patterns as well as natural linguistic ones—paired with a soft, quantitative simplicity metric that captures learning dynamics or functional pressures on language. Learning models with explicit simplicity biases exemplify this approach (for example, Hsu et al.,, 2011; Perfors et al.,, 2013; Rasin et al.,, 2021; Lan et al.,, 2022).
@@
This is not to say that inductive biases are no longer important in linguistics and language learning, nor that investigating inductive biases of humans and neural networks is not important: we will have much more to say on this in Section 4.4. Far from demoting inductive bias as a concern, language models open up the range of (possibly innate) inductive biases that we might look for in humans. The main point for linguistic theory is not to demote the importance of the learning problem, but to reshape it: away from explanation by constrained description, and toward a broader landscape of approaches and hypotheses.
3.2.4 The question of data quantity
It is still the case that, whatever language models learn about language, they do it using orders of magnitude more linguistic input data than human children are exposed to (Yedetore et al.,, 2023; Warstadt et al.,, 2023). Moreover, the learning trajectories of models and humans continue to show systematic differences (Chang and Bergen,, 2022; Evanson et al.,, 2023; Constantinescu et al.,, 2025). Taken together, these results suggest meaningful differences between learning in models and humans, both in terms of data requirements and patterns of learning. A form of the Poverty of the Stimulus argument is still alive in the form of the observation that even if neural networks acquire linguistic structure, they do not do so on the basis of the same amount and kind of data that a human child learner gets (Lan et al.,, 2024).
As reviewed above, it has been suggested that formal linguistic structure provides the key inductive bias that makes learning from small amounts of data possible. However, there is no evidence that linguistic structure is the major missing ingredient to close the gap between models and human children. [^12] Indeed, the analysis by Mollica and Piantadosi, (2019) suggests that syntactic structure makes up only a very small portion of the information necessary to learn a language. There have been many attempts to inject linguistic structure into neural models in various ways; some of the earliest successful deep learning approaches to language understanding were based on architectures that performed recursive computation in hierarchical parse trees (Socher et al.,, 2011, 2013). These approaches do succeed in promoting the kinds of structural generalizations that humans find intuitive, with the result that language learning from data is somewhat more sample efficient (Dyer et al.,, 2016; Futrell et al., 2019b, ; Wilcox et al., 2019b, ; Kim et al.,, 2019; Papadimitriou and Jurafsky,, 2020, 2023; Nandi et al.,, 2025). However, in terms of absolute performance, these approaches are still far from closing the gap between humans and neural networks.
It is entirely possible that the key to faster and more human-like learning is not particular built-in formal constraints, but rather more flexible architectures (such as Kolmogorov–Arnold Networks: Liu et al.,, 2024), or biases towards domain-general compositional reasoning (McCoy and Griffiths,, 2023; Yang and Piantadosi,, 2022), or different training regimes (Murty et al.,, 2023), or the incorporation of multimodal data which provides rich side information about the structure of the environment that is being described in language (Wang et al.,, 2023), or domain-general bounded-rational approaches to generalization such as the Tolerance Principle (Belth et al.,, 2021; Payne et al.,, 2021; Kodner,, 2022).
Overall, we believe the shoe is on the other foot for those who believe that an innate bias toward a particular formal linguistic structure is the factor that enables human learning from small amounts of data. Formal linguistics has not presented an alternative model with the demonstrated practical language-learning capabilities of neural models. The direction of developments in machine learning suggests that the gap between human and machine learning is more likely to be closed not through built-in domain-specific formal restrictions, but rather through more powerful domain-general learning algorithms.
The difference between human and neural network learning also raises the possibility that humans and neural networks are just so different that one is not informative about the other (Kodner et al.,, 2023). We believe not. First, as argued above, the success of neural networks weakens logical arguments that language cannot be learned without domain-specific formal constraints on language, and in general changes how we think about the role of formal restrictions in learning. Second, even if standard neural network training methods are not able to acquire linguistic structure on the basis of developmentally realistic data, the representations that they do eventually acquire based on more data are still informative about how language might be represented and processed in the brain (more in Section 4), even if the networks do not arrive at these representations along the same trajectory as the human child.
@@ 3.3 Language models and linguistic traditions
While language models may seem to involve concepts that are foreign to linguistics, they emerged in part from certain intellectual traditions of the study of language—although these traditions are not necessarily the ones that have been dominant in linguistics departments. Below we discuss the relationship between schools of linguistics and the development of language models and related language technologies.
@@ 3.3.1 The generative tradition of linguistics
A historical parallel is informative. In the early 19th century, while Newtonian mechanics provided a strong basis for parts of mechanical engineering, it did not on its own seem to provide a theory that answered pressing questions about the increasingly complex machines emerging from the Industrial Revolution, in particular steam engines. Carnot, (1824) developed an effective theory of such engines using a new ad-hoc concept of ‘moment of activity’, which eventually developed into the idea of entropy (Clausius,, 1865). This concept had to be discovered through engineering because the focus of purely theoretical physics was on understanding fundamental mechanics through mathematical methods of increasing elegance. Explaining the behavior of complex machines—which behavior certainly must reduce to more basic mechanics anyway—was not on the agenda. And yet from the practical project of understanding steam engines emerged a family of concepts that some theoretical physicists now view as more fundamental than even matter itself (Wheeler,, 1989).
@@
Turning back to linguistics: the reason that generative linguistic theory did not turn out to be helpful along the path toward language models was due to a similar narrowness in focus, which led it away from considering complex systems for dealing with language.
@@
Language models are, at bottom, models of the stream of language that is produced and comprehended by humans. The relevant theory is the theory of language use, production, comprehension, learning, and of human cognition more generally. Generative linguistics, on the other hand, has not focused on how to build a machine that produces and comprehends language, but rather how to build a language, conceived of as an abstract mental structure that gives rise to a mapping between meaning (that is, a logical form or conceptual–intensional representation) and form (that is, a phonological form or sensorimotor representation) (Chomsky,, 1995, 2005; Adger,, 2003; Hornstein et al.,, 2005). [^13] Early generative linguistics went further in its goals, claiming that the structures identified through generative analysis would not only provide a characterization of language, but also be a necessary and explicit part of any theory of how language could be learned and processed (Chomsky,, 1965, 1971; Bresnan,, 1982). See Stabler, (1983) for discussion of what he sees as confusion within the field as to whether or not generative linguistic theories are intended to be theories of the representations used by the brain during processing, as opposed to theories that constrain possible language. He concludes that Chomsky and others often conflated theories of grammar and theories of mental representation and processing, but that this position was not well grounded and that “it seems unlikely that linguists and psychologists really want to claim any such thing”. Furthermore, it was claimed that this kind of analysis must take center stage in the science of language, preceding any analysis of more complex systems for language processing, use, or learning, since these systems must operate in ways that make reference to the abstract structures of language (Chomsky,, 1965, Ch. 1).
@@
Given this focus, generative linguistics has had a huge influence on the engineering field that works on how to build languages: programming language design. Indeed, the basics of all modern programming languages are based on formal linguistic theory, incorporating explicit (often context-free) grammars and parsers, as a way of linking a stream of text to a representation of a computation. Programming languages are, perhaps, exactly the ‘esoteric dialects’ spoken by ‘elaborate machines’ in the Nabokov quote that starts this section.
@@
While this narrowness of focus was useful on its own terms, it has limits. There are potentially fundamental insights to be gained now from the analysis of messy, complex, practical systems, just as happened in physics, and as is now happening in linguistics.